If the greatest and the least resultants of two forces acting at a point at an angle x, then find the magnitude and direction of their resultant.

Grade:12th pass

1 Answers

Vikas TU
4 years ago
Dear student 
Let us understand by this example. 
Let the two forces be x and y
Greatest resultant ( parallel cosz=1) = x+y = 10
Least Resultant ( antiparallel cosz= -1) = x-y = 6
solving this to equation x =8N, y =2N
Each force is increased by 3N
Therefore x = 11N, y = 5N
Let angle be 90 degree . 
Resultant vector = ( x^2 + y^2 +2xycos90*) ^1/2
= (121 + 25)^1/2
= 146^1/2
= 12.08N
Good Luck